OpenStudy (anonymous):

the equation is 4r[2]-2r=0 [2] Means square root in the equation

OpenStudy (anonymous):

r=0

OpenStudy (anonymous):

\(4r^2-2r=0\), take 2r as a common factor, you get \(2r(2r-1)=0\). So r is either \(0\) or \(1/2\).

OpenStudy (anonymous):

[2] is square root.

OpenStudy (anonymous):

Is the question, \[4r \sqrt{2} - 2r =0\]

OpenStudy (anonymous):

Yes that is the question

OpenStudy (anonymous):

Whoops, sorry for the misunderstanding :D

OpenStudy (anonymous):

It's ok

OpenStudy (anonymous):

So if you take out r, the only possibility is r=0

OpenStudy (anonymous):

Thank you very much for your help!!

OpenStudy (anonymous):

welcome :)